HIGH – RESOLUTION LASER SPECTROSCOPY OF THE A^{3}Π_{1} ← X^{1}Σ^{+} SYSTEM OF ICl IN 0.8 μm REGION.

\underline{N. NISHIMIYA} and T. YUKIYA, Department of System Electronics and Information Technology, Department of Applied Computer Science, Tokyo Plolytechnic University, Iiyama 1583, Atsugi City, 243-0297 Kanagawa, JAPAN.

The doppler limited electronic vib–rotational absorption spectrum of A^{3}Π_{1} ← X^{1}Σ^{+} system in I^{35/37}Cl is measured in the 11300 – 13200 cm^{−1} region using a Ti:Sapphire Ring Laser (Coherent 899-21). The 56 bands belonging to (v′ = 0 ∼ 7) ← (v′′ = 0 ∼ 6) in I^{35/37}Cl are assigned. The Dunham coefficients of the X^{1}Σ^{+} state are calculated based on a global reduced least squares fitting procedure. The Y′′_{10} and Y′′_{01} reported by Hedderich et al.^{a} are taken into the calculation because we do not have sufficient spectrum belonging to v′′= 0 ∼ 2. In the A^{3}Π_{1} state, the T′_{v}, B′_{v}_{e/f}, D′_{v} and H′_{v} are determined instead of the Dunham coefficients to avoid the distortion of the potential. To investigate the accuracy of Dunham coefficients determined in the X^{1}Σ^{+} state, the absorption lines of 4′ ← 0′′ and 5′ ← 0′′ bands are assigned in the 14438 – 14491 cm^{−1} region where measurements are performed intermittently. Those lines are well positioned within the limit of the standard deviation.

aH. G. Hedderich and P. F. Bernath, J. Mol. Spectrosc. 155,384–392(1992)